OpenStudy (anonymous):

integral from 0 to 1 of 8x/((4x^2 + 1)^2) using substitution? I understand that u should = 4x^2 + 1 Then change the limits for u to 1 to 5 by filling the original limits in for u. And du = 8x dx. So I'm left with u^-2. What now?

OpenStudy (anonymous):

find the anti derivative using the power rule backwards

OpenStudy (anonymous):

Okay antiderivative is 1/2 x^-1 then

OpenStudy (anonymous):

\[\int u^ndu=\frac{u^{n+1}}{n+1}\] in your case \(n=-2\)

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

good question. Atleast you did not ask "AY YO I HAVE A QUESTION"

OpenStudy (anonymous):

u= 4x^2+1

OpenStudy (anonymous):

should be \[\int u^{-2}du=-u^{-1}=-\frac{1}{u}\]

OpenStudy (anonymous):

evaluate at 5, 1, and subtract

OpenStudy (jhannybean):

Are you concurrently taking Calc 3 and Calc 4 together?

OpenStudy (anonymous):

wait no that's not right is it. Okay so it's actually - 1/u. Okay that makes sense. Then I fill in the new limits and subtract?

OpenStudy (anonymous):

calc 4? is that before calc 5, 6, and 7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright that works. Thanks satellite73